OpenStudy (anonymous):

Differentiate this: y= (3-x)^6

OpenStudy (freckles):

you could use chain rule

OpenStudy (freckles):

\[\frac{d}{dz} (z^6)_{z=3-x} \cdot (3-x)'\]

OpenStudy (anonymous):

like this: 6(3-x)^5 x (-1)

OpenStudy (anonymous):

?

OpenStudy (freckles):

is that one x suppose to be a multiplying sign?

OpenStudy (freckles):

\[6(3-x)^{5} \cdot (-1)\]

OpenStudy (anonymous):

Yes. So it would be -6(3-zx^5

OpenStudy (anonymous):

-6(3 - x)^5

OpenStudy (freckles):

yep